Ghose and Han​ (2014) found that the elasticity of demand for Google Play apps is −3.7. This elasticity applies to a small college town where approximately​ 1,000 apps per month are sold. If price rises by 10​%, what would be the effect on quantity​ demanded? The quantity demanded will decrease by 37 percent. Would revenue rise or​ fall? Revenue would fall .

What is the percentage change in​ revenue using information above? ​(Enter your response rounded to two decimal​ places.)
Answer #1

Let initial price and quantity be P0 and Q0 respectively.

Initial revenue (R0) = P0 x Q0

After price rises,

New price (P1) = P0 x 1.10

New quantity (Q1) = Q0 x (1 - 0.37) = Q0 x 0.63

New revenue (R1) = P1 x Q1 (P0 x 1.1) x (Q0 x 0.63) = (1.1 x 0.63) x (P0 x Q0)= 0.693 x R0

% Change in revenue = (R1 - R0)/R0 = [(0.693 x R0) - R0] / R0 = (0.693 - 1) / 1 = -0.3070 = -30.70%